OpenStudy (anonymous):

Can you please explain (in terms of particles and energy) what would happen to a blown-up party balloon if it is placed in the refrigerator ?

OpenStudy (aaronq):

when you inflate a balloon and room temp, the particles moving about have a certain kinetic energy. Collisions between the particles with the inner surface of the balloon exert pressure on the surface. When you lower the temperature of the environment (put the balloon in a fridge) you lower the kinetic energy of the particles moving inside the balloon. As a result, the particles have less frequent and less forceful collision withe inner surface of the balloon. The balloon shrinks. This result can be explored by using the ideal gas law: \(\sf PV=nRT\) Assume that P= 1 atm, and V = 1 L at room temp, 298 K \(\sf \dfrac{P_1V_1}{nRT_1}=\dfrac{P_2V_2}{nRT_2}\) because the external pressure, P, the number of particles, n, and the constant R dont change, we can exclude them from the equation. Say that the temp in the fridge is 20 degrees lower than the room temp. \(\sf \dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\rightarrow \dfrac{1~L}{298~K}=\dfrac{V_2}{278~K} \rightarrow V_2=0.933~L\) The volume decreased from 1 L to 0.933 L by dropping the temp by 20 degrees.